# VCR: Visual Caption Restoration

### v2

@@ Abstract
We introduce Visual Caption Restoration (VCR), a novel vision-language task that challenges models to accurately restore partially obscured texts using pixel-level hints within images. This task stems from the observation that text embedded in images is intrinsically different from common visual elements and natural language due to the need to align the modalities of vision, text, and text embedded in images. While numerous works have integrated text embedded in images into visual question-answering tasks, approaches to these tasks generally rely on optical character recognition or masked language modeling, thus reducing the task to mainly text-based processing. However, text-based processing becomes ineffective in VCR as accurate text restoration depends on the combined information from provided images, context, and subtle cues from the tiny exposed areas of masked texts. We develop a pipeline to generate synthetic images for the VCR task using image-caption pairs, with adjustable caption visibility to control the task difficulty. With this pipeline, we construct a dataset for VCR called VCR-wiki using images with captions from Wikipedia, comprising English and Chinese entities in both easy and hard configurations. Our results reveal that current vision language models significantly lag behind human performance in the VCR task, and merely fine-tuning the models on our dataset does not lead to notable improvements. We release VCR-wiki and the data construction code to facilitate future research.
@@ 1 Introduction
Recent advances in large language models, such as ChatGPT [51, 50] and Llama [62], have spurred significant interest and progress in the field of vision-language models. With models like GPT-4V [50] and Llava [38, 39, 40] blending textual and visual information, the intersection of computer vision and natural language processing has become a vibrant research frontier. These integrated models aim to leverage the potential of vision and language modalities to understand and interpret multimedia content more effectively.
@@
Amidst this evolving landscape, we introduce Visual Caption Restoration (VCR), a novel vision-language task designed to challenge existing models uniquely. VCR challenges these models to restore obscured texts within images, a task that demands an intricate synthesis of text, vision, and text embedded in the image. The VCR task is grounded in two key insights: (1) text embedded within images, with its characteristics different from common visual elements, represents a distinct modality that requires careful alignment of vision, textual data, and the structure of written texts, and (2) neuroscience findings that suggest that humans are proficient in recognizing partially occluded objects through sophisticated visual and cognitive processes [61, 52, 63, 16, 34]. By leveraging these insights, VCR seeks to explore how well vision-language models can handle texts embedded within images, aligning visual elements and natural language to mimic human-like multimodal understanding and recognition.
@@
The Visual Question Answering (VQA) task [3, 65, 47, 56] has been a popular benchmark in assessing how well models align and interpret visual and linguistic information. Traditional VQA approaches, however, predominantly focus on direct queries about visible elements in images and do not address the nuanced relationship between textual content embedded within the image and the overall image context. This gap underscores the limited capabilities of current models in processing integrated visual-textual data, particularly when the textual component, which plays a critical role, is partially obscured or altered.
@@
To address these limitations, our VCR task builds on the premise that effective text restoration from images requires an integrated understanding beyond the capabilities of current VQA benchmarks. For example, in extreme cases, models rely on existing Optical Character Recognition (OCR) system to extract text from documents [56, 7]. The extracted text is then used as context for generating answers, without a true semantic alignment between the text and the visual elements of the document. This approach, while effective in simple scenarios, falls short in more complex settings where text is intricately woven into the visual narrative of the image.
@@
To develop the VCR task, in this work, we introduce a pipeline for generating synthetic images that allows for manipulation of the visibility of the textual components of the image. This not only enhances the challenge posed by the task, but also provides a scalable way to adjust task difficulty. The resulting dataset, VCR-wiki, comprises 2.11M English data and 346K Chinese data sourced from Wikipedia, featuring captions in both languages across ‘easy’ and ‘hard’ difficulty levels. Our evaluations indicate that existing vision-language models significantly underperform compared to human benchmarks, underscoring the need for novel model architectures and training paradigms specifically geared towards this complex intermodal alignment.
@@
Introduce the Visual Caption Restoration (VCR) task to challenge vision-language models to restore occluded texts in images.
@@
Develop a pipeline for generating synthetic images with embedded text that allows for adjusting visibility of such text, thus providing a rich testing environment for VCR.
@@ 2 VCR Task Description
For better clarity, we define text embedded in image () as text incorporated within the image. The term visual image (VI) pertains to the portion of the image that excludes the text embedded in the image. The string text (ST) is not part of the image itself, but is associated with it as a distinct textual element. It is usually the question prompt in the form of natural language, for example, ‘What are the covered texts in the image? Please only guess the covered texts without outputting an explanation.’. Consequently, an element of a Visual Caption Restoration (VCR) task can be expressed as , where is represented as a string and both and are presented in image form. This notation does not imply that and can be physically separated into two distinct image components. Instead, this definition is adopted merely to facilitate a clearer explanation of the concepts involved. Please refer to Figure 2 for an illustration of , , and .
@@
The first aspect of importance stems from discoveries in neuroscience about human cognitive abilities to recognize partially occluded objects [16, 34]. Although existing models can recognize objects and texts in images, they often struggle with the complexity of occluded objects due to significant information loss in the images. In contrast, humans excel at filling in missing information using a combination of low-level visual processing and high-level cognitive functions, such as those managed by the prefrontal cortex. This cortical area is known to handle complex cognitive processes such as decision making and memory retention, which are essential for integrating fragmented visual input into coherent objects. We believe that the occlusion restoration task serves as a probe that can effectively distinguish low-level recognition and high-level cognition involving reasoning. In addition, understanding these neural mechanisms can inspire new algorithms capable of mimicking human-like perception and interpretation in dynamic, real-world conditions where occlusion is common.
The second aspect underscores the unique challenge presented by the VCR task, distinguishing it significantly from existing benchmarks, such as traditional VQA or the occluded object restoration task. By occluding texts instead of common visual objects, VCR targets the models’ text-image alignment capability, which is one of the major challenges for vision-language models. VCR mandates that models align textual and visual information in a manner that replicates human-like understanding involving the utilization of both textual and visual clues. This task requires a deep integration of visual (), embedded textual (), and contextual interpretation across modalities, pushing beyond simple text extraction as performed in OCR tasks. In OCR, the focus is primarily on recognizing visible characters, often without the need to understand their contextual relevance within the image narrative. In contrast, VCR introduces complexity by requiring the model to use available partial texts and the visual context collaboratively to reconstruct the obscured content accurately. This not only tests the model’s ability to process - modalities effectively, but also challenges it to maintain internal consistency, akin to human cognitive processes where context and visual clues guide understanding and response. Besides, the difficulty of the task can be adjusted by altering the extent of text occlusion, offering a scalable and flexible framework for systematically enhancing model capabilities in text-visual alignment and semantic comprehension. This rigorous testing ground will help evolve vision-language models to better grasp the nuanced interplay between text and imagery.
@@ 3 Dataset Creation
Within , we use spaCy to randomly select several 5-grams in the caption for masking. To ensure the restoration process is doable by a human without too much professional domain knowledge, the 5-grams do not contain numbers, person names, religious or political groups, facilities, organizations, locations, dates and time labeled by spaCy, and the total masked token does not exceed 50% of the tokens in the caption. We exclude all instances that do not have a single eligible 5-gram. The selected 5-grams are partially obscured by a white rectangle that reveals only the upper and lower parts of the text, with the proportion of coverage varying to adjust task difficulty. Furthermore, to assess the impact of on model performance, we create an ablation for each image, maintaining the resolution of the + image, but retaining only the part in the center of the image.
@@ 4.1 Models
We evaluate several most advanced proprietary models with their provided APIs. The evaluated models include GPT-4o (gpt-4o-2024-0513), GPT-4 Turbo (gpt-4-turbo-2024-04-09), GPT-4V (gpt-4-1106-vision-preview) [51, 50], Claude 3 Opus (claude-3-opus-20240229), Claude 3.5 Sonnet (claude-3-5-sonnet-20240620) [2], Gemini 1.5 pro (gemini-1.5-pro-001) [59], Reka Core (reka-core-20240501) [60], and Qwen-VL-Max (tested on May 2024) [4].
@@
We evaluate open-source models with the best performance on the OpenVLM Leaderboard222We selected the highest-performing open-source models with fewer than 40 billion parameters from the OpenVLM Leaderboard as of May 23, 2024. Details are available at https://huggingface.co/spaces/opencompass/open_vlm_leaderboard. and state-of-the-art Chinese VLM models. The evaluated models include InternVL-Chat-V1.5[11], MiniCPM-Llama3-V2.5 [25], InternLM-XComposer2-VL-7B [14], CogVLM2-Llama3-19B-Chat [67], Idefics2-8B [32], Yi-VL-34B [1], Yi-VL-6B [1], Qwen-VL-Chat [4], DeepSeek-VL-7B-Chat [43], DeepSeek-VL-1.3B-Chat [43], Monkey [41, 35] and DocOwl-1.5 [22]. Out of these models, Idefics2-8B is an English-only model, and CogVLM2-Llama3-19B-Chat has its Chinese variant, CogVLM2-Llama3-19B-Chinese-Chat. Please refer to Table 2 for the model specifications.
@@
To test whether VLMs can learn to conduct VCR via fine-tuning, we select two models from the open-sourced models, CogVLM2-Llama3-19B-Chat and MiniCPM-Llama3-V2.5, and fine-tune them on a subset of VCR’s training set.
More specifically, we fine-tune CogVLM2-Llama3-19B-Chat and MiniCPM-Llama3-V2.5 in the English Hard configuration, and CogVLM2-Llama3-19B-Chinese-Chat and MiniCPM-Llama3-V2.5 on the Chinese Hard configuration. The models are finetuned using LoRA [23] with and . We adopt the schedule-free AdamW optimizer [12] with a learning rate . The effective batch size is 64. Each model is trained on the first 16,000 examples of the training set for 1 epoch. All fine-tuning experiments are performed on a single node with 4 NVIDIA L40S 48G GPUs.
@@ 4.2 Metrics
We report the two metrics below in our experiment section to measure the restoration quality: Exact Match (), which measures whether the restored -gram totally matches the ground-truth ; and Jaccard Index (), which measures the similarity of and as bag-of-words.
Exact Match (), which measures whether the restored -gram totally matches the ground-truth ;
Jaccard Index (), which is a more relaxed metric that measures the similarity of and as bag-of-words.
where represents the set of tokens in .
@@ 4.3 Experimental Results
Please refer to the exact match score and the Jaccard-index of the evaluation in Table 4.3.
Open-Source Models.
We evaluate each open-source model based on the whole 5,000 examples in the test set. Note that Idefics2-8B only supports the English task, hence it has no evaluation score on the Chinese task.
Although achieving state-of-the-art performance on the Open VLM leaderboard, almost all the tested models achieve a low exact match accuracy in the English Easy configuration and fail on the other settings. The best open-source model across the 4 configurations (English Easy, English Hard, Chinese Easy, and Chinese Hard) is CogVLM2-Llama3-Chat. This might be attributed to its pretraining process and the special architecture. We also notice that VI has a negative impact for most models on the exact match scores (), which means that the image information is not properly utilized. The best performed open-source model, CogVLM2-Llama3-Chat, and its fine-tuned version have positive , except for the Chinese Hard configuration. This indicates that information from VI could help improve the model performance on VCR.
For different languages, we noticed a large performance drop when testing the model in Chinese configurations, despite the fact that all models claim to have basic English-Chinese duolingual capabilities. This is somehow surprising, since Chinese characters, due to their logographic nature, may exhibit a higher degree of recognizability compared to languages that use alphabetic scripts in one order [69, 79].
@|c|c|p4cmC2cmC2.3cmC2.3cmC2cmC2.3cmC2.3cmC2cm|@
@@
English Easy Closed-source models
Claude 3 Opus - -15 -10.74
@@
Idefics2 8B -12.02 -19.03
InternLM-XComposer2-VL 7B 0.24 -1.14
InternVL-V1.5 25.5B -60.41 -35.68
MiniCPM-V2.5 8B -8.25 -9.96
MiniCPM-V2.5-FT 8B -3.67 -3.22
@@
Hard Closed-source models
Claude 3 Opus - -12.2 -12.48
@@
Idefics2 8B -0.29 -2.64
InternLM-XComposer2-VL 7B -0.22 -0.72
InternVL-V1.5 25.5B -4.5 -9.67
MiniCPM-V2.5 8B -0.55 -1.43
MiniCPM-V2.5-FT 8B 0.12 0.38
@@
Chinese Easy Closed-source models
Claude 3 Opus - -0.1 1.49
@@
Open-source models
CogVLM2-Chinese 19B 2.54 3.91
CogVLM2-Chinese-FT 19B 1.84 1.02
@@
Hard Closed-source models
Claude 3 Opus - 0.2 1.13
@@
Open-source models
CogVLM2-Chinese 19B -1.32 -2.16
CogVLM2-Chinese-FT 19B -3.51 -3.61
@@
Moreover, we found that models, such as internlm-xcomposer2, are good at OCR and understanding image documents (as demonstrated by DocOwl 1.5 and Monkey) still have the potential to be improved in the VCR task. This highlights the unique and indispensable role of VCR in the current suite of benchmarks. Excelling in other document-related benchmarks does not guarantee similar performance in VCR tasks, emphasizing VCR’s distinct challenges and value.
Closed-Source Models.
We evaluate every closed-source model with the first 500 examples in the test set. In English tasks, GPT-4o scores the best among the models that have not been finetuned. Even though GPT-4 series support Chinese, we found that GPT-4V (gpt-4-1106-vision-preview) is not able to recognize Chinese characters embedded in the image even without any occlusion. Thus, we do not test GPT-4V on Chinese tasks.
In English configurations, closed-source models outperform all open-source models except CogVLM2, which indicates that model scaling might help improve performance on the VCR task. However, compared with the human evaluation results in Section 4.4, we notice a large performance gap, especially in the English Hard configuration. This shows significant room for improvement in the current state-of-the-art models.
Refer to Table A to compare open and closed source models using the same 500 test cases.

### v3

@@ Abstract
We introduce Visual Caption Restoration (VCR), a novel vision-language task that challenges models to accurately restore partially obscured texts using pixel-level hints within images. This task stems from the observation that text embedded in images intrinsically differs from common visual elements and natural language due to the need to align the modalities of vision, text, and text embedded in images. While numerous works have integrated text embedded in images into visual question-answering tasks, approaches to these tasks generally rely on optical character recognition or masked language modeling, thus reducing the task to mainly text-based processing. However, text-based processing becomes ineffective in VCR as accurate text restoration depends on the combined information from provided images, context, and subtle cues from the tiny exposed areas of masked texts. We develop a pipeline to generate synthetic images for the VCR task using image-caption pairs, with adjustable caption visibility to control the task difficulty. With this pipeline, we construct a dataset for VCR called VCR-wiki using images with captions from Wikipedia, comprising English and Chinese entities in both easy and hard configurations. Our results reveal that current vision language models significantly lag behind human performance in the VCR task, and merely fine-tuning the models on our dataset does not lead to notable improvements. We release VCR-wiki and the data construction code to facilitate future research.
@@ 1 Introduction
Recent advances in large language models, such as ChatGPT (Ouyang et al., 2022; OpenAI et al., 2023) and Llama (Touvron et al., 2023), have spurred significant interest and progress in the field of vision-language models. With models like GPT-4V (OpenAI et al., 2023) and LLaVA (Liu et al., 2023a; 2024a; b) blending textual and visual information, the intersection of computer vision and natural language processing has become a vibrant research frontier. These integrated models aim to leverage the potential of vision and language modalities to understand and interpret multimedia content more effectively.
@@
Amidst this evolving landscape, we introduce VCR, a novel vision-language task designed to challenge existing models uniquely. VCR challenges these models to restore obscured texts within images, which demands an intricate synthesis of text, vision, and text embedded in the image. The VCR task is grounded in two key insights: (1) text embedded within images, with its characteristics different from common visual elements, represents a distinct modality that requires careful alignment of vision, textual data, and the structure of written texts, and (2) neuroscience findings that suggest that humans are proficient in recognizing partially occluded objects through sophisticated visual and cognitive processes (Thinés et al., 2013; Pessoa et al., 1998; van Lier & Gerbino, 2015; Fyall et al., 2017; Li et al., 2023a). By leveraging these insights, VCR seeks to explore how well vision-language models can handle texts embedded within images, aligning visual elements and natural language to mimic human-like multimodal understanding and recognition.
@@
The Visual Question Answering (VQA) task (Antol et al., 2015; Wang et al., 2018; Mishra et al., 2019b; Singh et al., 2019) has been a popular benchmark in assessing how well models align and interpret visual and linguistic information. Traditional VQA approaches, however, predominantly focus on direct queries about visible elements in images and do not address the nuanced relationship between textual content embedded within the image and the overall image context. This gap underscores the limited capabilities of current models in processing integrated visual-textual data, particularly when the textual component, which plays a critical role, is partially obscured or altered.
@@
To address these limitations, our VCR task builds on the premise that effective text restoration from images requires an integrated understanding beyond the capabilities of current VQA benchmarks. For example, in extreme cases, models rely on existing Optical Character Recognition (OCR) system to extract text from documents (Singh et al., 2019; Borisyuk et al., 2018). The extracted text is then used as context for generating answers without a true semantic alignment between the text and the visual elements of the document. This approach, while effective in simple scenarios, falls short in more complex settings where text is intricately woven into the visual narrative of the image.
@@
To develop the VCR task, in this work, we introduce a pipeline for generating synthetic images that allows for manipulation of the visibility of the textual components of the image. This not only enhances the challenge posed by the task but also provides a scalable way to adjust task difficulty. The resulting dataset, VCR-wiki, comprises 2.11M English data and 346K Chinese data sourced from Wikipedia, featuring captions in both languages across ‘easy’ and ‘hard’ difficulty levels. Our evaluations indicate that existing vision-language models significantly underperform compared to human benchmarks, underscoring the need for novel model architectures and training paradigms specifically geared towards this complex intermodal alignment.
@@
Introduce the VCR task to challenge vision-language models to restore occluded texts in images.
@@
Develop a pipeline for generating synthetic images with embedded text that allows for adjusting the visibility of such text, thus providing a rich testing environment for VCR.
@@ 2 VCR Task Description
For better clarity, we define text embedded in image () as text incorporated within the image. The term visual image (VI) pertains to the portion of the image that excludes the text embedded in the image. The string text (ST) is not part of the image itself but is associated with it as a distinct textual element. It is usually the question prompt in the form of natural language, for example, ‘What are the covered texts in the image? Please only guess the covered texts without outputting an explanation.’. Consequently, an element of a VCR task can be expressed as , where is represented as a string and both and are presented in image form. This notation does not imply that and can be physically separated into two distinct image components. Instead, this definition is adopted merely to facilitate a clearer explanation of the concepts involved. Please refer to Figure 3 for an illustration of , , and .
@@
The first aspect of importance stems from discoveries in neuroscience about human cognitive abilities to recognize partially occluded objects (Fyall et al., 2017; Li et al., 2023a). Although existing models can recognize objects and texts in images, they often struggle with the complexity of occluded objects due to significant information loss in the images. In contrast, humans excel at filling in missing information using a combination of low-level visual processing and high-level cognitive functions, such as those managed by the prefrontal cortex. This cortical area is known to handle complex cognitive processes such as decision making and memory retention, which are essential for integrating fragmented visual input into coherent objects. We believe that the occlusion restoration task serves as a probe that can effectively distinguish low-level recognition and high-level cognition involving reasoning. In addition, understanding these neural mechanisms can inspire new algorithms capable of mimicking human-like perception and interpretation in dynamic, real-world conditions where occlusion is common.
The second aspect underscores the unique challenge presented by the VCR task, distinguishing it significantly from existing benchmarks, such as traditional VQA or the occluded object restoration task. By occluding texts instead of common visual objects, VCR targets the models’ text-image alignment capability, which is one of the major challenges for vision-language models. VCR mandates that models align textual and visual information in a manner that replicates human-like understanding involving the utilization of both textual and visual clues. This task requires a deep integration of visual (), embedded textual (), and contextual interpretation across modalities, pushing beyond simple text extraction as performed in OCR tasks. In OCR, the focus is primarily on recognizing visible characters, often without the need to understand their contextual relevance within the image narrative. In contrast, VCR introduces complexity by requiring the model to use available partial texts and the visual context collaboratively to reconstruct the obscured content accurately. This not only tests the model’s ability to process - modalities effectively but also challenges it to maintain internal consistency, akin to human cognitive processes where context and visual clues guide understanding and response. Besides, the difficulty of the task can be adjusted by altering the extent of text occlusion, offering a scalable and flexible framework for systematically enhancing model capabilities in text-visual alignment and semantic comprehension. This rigorous testing ground will help evolve vision-language models to better grasp the nuanced interplay between text and imagery. We show an example of how humans would solve the VCR task in ‘hard’ difficulty in Figure 2.
@@ 3 Dataset Creation
Within , we use spaCy to randomly select several 5-grams in the caption for masking. To ensure the restoration process is doable by a human without too much professional domain knowledge, the 5-grams do not contain numbers, person names, religious or political groups, facilities, organizations, locations, dates, and time labeled by spaCy, and the total masked token does not exceed 50% of the tokens in the caption. We exclude all instances that do not have a single eligible 5-gram. The selected 5-grams are partially obscured by a white rectangle that reveals only the upper and lower parts of the text, with the proportion of coverage varying to adjust task difficulty. Furthermore, to assess the impact of on model performance, we create an ablation for each image, maintaining the resolution of the + image, but retaining only the part in the center of the image.
@@ 4.1 Models
We evaluate several most advanced proprietary models with their provided APIs. The evaluated models include GPT-4o (gpt-4o-2024-0513), GPT-4 Turbo (gpt-4-turbo-2024-04-09), GPT-4V (gpt-4-1106-vision-preview) (Ouyang et al., 2022; OpenAI et al., 2023), Claude 3 Opus (claude-3-opus-20240229), Claude 3.5 Sonnet (claude-3-5-sonnet-20240620) (Anthropic, 2024), Gemini 1.5 pro (gemini-1.5-pro-001) (Team et al., 2024a), Reka Core (reka-core-20240501) (Team et al., 2024b), and Qwen-VL-Max (tested in May 2024) (Bai et al., 2023).
@@
We evaluate open-source models with the best performance on the OpenVLM Leaderboard222We selected the highest-performing open-source models with fewer than 40 billion parameters from the OpenVLM Leaderboard as of May 23, 2024 and their later versions. Details are available at https://huggingface.co/spaces/opencompass/open_vlm_leaderboard. and state-of-the-art Chinese VLM models. The evaluated models include Cambrian-1 (Tong et al., 2024), CogVLM2-19B (Hong et al., 2024), DeepSeek-VL-7B-Chat (Lu et al., 2024), DocOwl-1.5-Omni (Hu et al., 2024a), Idefics3-8B (Laurençon et al., 2024), InternLM-XComposer2-VL-7B (Dong et al., 2024a), InternLM-XComposer2.5-VL (Zhang et al., 2024), InternLM-XComposer2-VL-4K (Dong et al., 2024b), Llama-3.2 (Dubey et al., 2024), MiniCPM-V2.5 (Hu et al., 2024b), Monkey (Liu et al., 2024b; Li et al., 2023b), Pixtral (Mistral, 2024), Qwen-VL-Chat (Bai et al., 2023), Qwen2-VL (Wang et al., 2024), and Yi-VL (01.AI et al., 2024). Out of these models, Cambrian-1, Idefics3 and Llama-3.2 are English-only models, and CogVLM2-Llama3-19B-Chat has its Chinese variant, CogVLM2-Llama3-19B-Chinese-Chat. Please refer to Table 6 for the model specifications.
@@
To test whether VLMs can learn to conduct VCR via fine-tuning, we select three models from the open-sourced models: CogVLM2-Llama3-19B-Chat, MiniCPM-Llama3-V2.5, and Qwen2-VL-7B-Instruct, and fine-tune them on a subset of VCR’s training set.
More specifically, we fine-tune CogVLM2-Llama3-19B-Chat, MiniCPM-Llama3-V2.5, and Qwen2-VL-7B-Instruct in the English Hard configuration, and CogVLM2-Llama3-19B-Chinese-Chat, MiniCPM-Llama3-V2.5, and Qwen2-VL-7B-Instruct on the Chinese Hard configuration. The models are finetuned using LoRA (Hu et al., 2022) with and . We adopt the schedule-free AdamW optimizer (Defazio et al., 2024) with a learning rate . The effective batch size is 64. Each model is trained on the first 16,000 examples of the training set for 1 epoch. All fine-tuning experiments are performed on a single node with 4 NVIDIA L40S 48G GPUs.
@@ 4.2 Metrics
We report the two metrics below in our experiment section to measure the restoration quality: , which measures whether the restored -gram totally matches the ground-truth ; and , which measures the similarity of and as bag-of-words.
4.3 Relationship to Other Benchmarks.
We evaluated 38 Vision-Language Models (VLMs) across 23 different benchmarks, using the VLM performance scores as features of each benchmark to compute a correlation matrix. Based on this matrix, in Figure 4, we applied K-Means clustering and visualized the results in 2D by plotting the first two principal components derived from the correlation matrix rows for each benchmark. Additionally,
we provide a heatmap of the correlation matrix in Appendix B.
and were excluded from these processes due to the limited availability of VLMs that support Chinese. From Figure 4, we observe that shows a tentative similarity to ChartQA and TextVQA, as all three benchmarks evaluate the ability to extract and reason about text from natural images and documents. However, does not exhibit significant similarity to the other benchmarks. On the other hand, stands apart from all other benchmarks. We attribute this to the fact that emphasizes caption recovery with minimal pixel-level information, a skill not tested by any of the other benchmarks. Therefore, we assert that the VCR series benchmarks assess unique aspects of VLMs that are not covered by any of the other benchmarks in our evaluation.
@@ 4.4 Experimental Results
Table 4.4 presents the exact match scores and Jaccard indices for our evaluations. Additionally, Figure 5 illustrates a comparative analysis of our fine-tuned models against their base counterparts across all four VCR settings. We also provide experiment results on smaller test sets of VCR-wiki for faster benchmarking. Please refer to Table A and A in the Appendix.
The VCR task and the VCR-wiki dataset are specifically designed to advance the development of future VLMs. In this section, we conduct a detailed analysis of the performance of current state-of-the-art (SOTA) models on the VCR task, highlighting key insights through comparative evaluations.
VCR Remains a Challenging Task for SOTA VLMs.
Despite observing high-performing models on the and settings, such as Qwen2-VL, the majority of recent models continue to struggle, particularly under the more demanding hard settings where performance metrics approach zero. This underscores the inherent difficulty of the VCR task and indicates that significant advancements are still required to achieve robust performance across all settings.
Enhanced OCR Capabilities Do Not Necessarily Translate to Improved VCR Performance.
Our analysis reveals that models proficient in OCR, such as InternLM-XComposer2-VL, and those excelling in image document understanding, like DocOwl 1.5 and Monkey, demonstrate subpar performance across most VCR settings. This discrepancy suggests that while these models can accurately recognize text within images, they lack the advanced reasoning capabilities required to effectively interpret and utilize this information within the context of the VCR task. This finding also highlights a fundamental distinction between OCR tasks and the more complex VCR task.
Language-Specific Performance: Need for Enhanced Multilingual Capabilities.
A significant performance degradation is observed when models are evaluated on Chinese configurations, despite assertions of basic English-Chinese bilingual capabilities. This decline is particularly surprising given the logographic nature of Chinese characters, which theoretically offer higher recognizability compared to alphabetic scripts (Wu et al., 2024; Zhao et al., 2022). These results indicate a critical need for targeted improvements in multilingual support to ensure consistent performance across different languages.
Model Size Does Not Guarantee Superior Performance.
Comparative analysis between Llama-3.2-11B and Llama-3.2-90B models reveals that both exhibit similar performance levels on the and settings. This observation suggests that merely increasing model size does not inherently enhance VCR performance. Instead, advancements in the cognitive abilities of models, achieved through improved training strategies, reasoning frameworks, or architectural innovations, are essential for meaningful performance gains in VCR tasks.
Model Resolution Is Not Directly Correlated with Performance Enhancement.
InternLM-XComposer2-VL-4K, despite its higher resolution, demonstrates significantly lower performance on the setting compared to its lower-resolution counterparts. This decline may result from more aggressive image partitioning strategies that disrupt the spatial continuity of text or from more intensive pixel or token compression techniques that lead to the loss of crucial local details. Both factors are critical for the accurate interpretation required in VCR tasks.
Inclusion of Input Images Negatively Impacts Performance.
The addition of generally results in negative performance changes (), indicating that the image information is not being effectively leveraged by the models. This negative impact may stem from the importance of key information locations, which could be compromised by image partitioning strategies that fail to preserve spatial relationships essential for accurate reasoning.
Model Architecture Influences Performance Gains from VCR-wiki Finetuning.
Finetuning on the VCR-wiki dataset yields varying performance improvements across different model architectures. Specifically: 1) CogVLM2 exhibits substantial performance enhancements across all four settings post-finetuning, suggesting that its architectural design is well-suited for the type of image-text reasoning required by the VCR task when provided with appropriate data and training. 2) MiniCPM-V2.5 shows only marginal performance increases from an already low baseline, particularly in the and settings. This limited improvement indicates potential architectural limitations that hinder its ability to effectively address the complexities of the VCR task. 3) Qwen2-VL-7B maintains relatively high performance both before and after finetuning, implying that the model is sufficiently pre-trained on relevant tasks to perform well on the VCR task without extensive additional training.
We hope that through controlled variables, the VCR-wiki dataset delivers fully comparable results across languages, difficulty levels, image inclusion, and fine-tuning stages. Each comparison is intended to guide specific and targeted improvements in VLM development.
@—c—c—c—p4.5cmC2cmC2.3cmC2.3cmC1.8cmC2.3cmC2.3cmC1.8cm—@
@@
English Easy Closed Claude 3 Opus - -15 -10.74
@@
Idefics3* 8B -5.44 -6.78
InternLM-XComposer2-VL 7B 0.24 -1.14
InternLM-XComposer2-VL-4K 7B 1.60 3.37
InternLM-XComposer2.5-VL* 7B 15.97 13.09
InternVL-V2* 40B -3.04 -2.47
Llama-3.2* 11B 12.32 9.47
Llama-3.2* 90B 9.48 5.59
MiniCPM-V2.5 8B -8.25 -9.96
@@
Pixtral* 12B 6.81 9.65
@@
Hard Closed Claude 3 Opus - -12.2 -12.48
@@
Idefics3* 8B 0.23 0.79
InternLM-XComposer2-VL 7B -0.22 -0.72
InternLM-XComposer2-VL-4K 7B 0.02 -0.00
InternLM-XComposer2.5-VL* 7B -0.18 -0.89
InternVL-V2* 40B -6.06 -7.71
Llama-3.2-Vision* 11B 7.17 8.90
Llama-3.2-Vision* 90B 1.85 1.00
MiniCPM-V2.5 8B -0.55 -1.43
@@
Pixtral* 12B -0.20 -0.46
@@
Chinese Easy Closed Claude 3 Opus - -0.1 1.49
@@
Open CogVLM2-Chinese 19B 2.54 3.91
@@
InternLM-XComposer2-VL 7B 0.04 0.04
InternLM-XComposer2-VL-4K 7B 0.00 -1.05
InternLM-XComposer2.5-VL* 7B -0.12 -2.01
InternVL-V2* 40B 4.84 9.69
MiniCPM-V2.5 8B -0.95 -4.9
@@
Hard Closed Claude 3 Opus - 0.2 1.13
@@
Open CogVLM2-Chinese 19B -1.32 -2.16
@@
InternLM-XComposer2-VL 7B -0.02 0.46
InternLM-XComposer2-VL-4K 7B 0.00 -0.04
InternLM-XComposer2.5-VL* 7B -0.01 -0.48
InternVL-V2* 40B -0.26 -0.74
MiniCPM-V2.5 8B 0.01 -0.5